OpenStudy (zombiekitty1296):

Solve for x: −4|x + 5| = −16

pooja195 (pooja195):

Hi there! Welcome to openstudy :-) −4|x + 5| = −16 Start by adding 4 to both sides −4+4|x + 5| = −16+4 -16+4=?

OpenStudy (zombiekitty1296):

-12?

pooja195 (pooja195):

Good |x + 5|=-12 absolute values have 2 solutions neg and ps so x + 5=-12 x + 5=12 solve both for x

OpenStudy (zombiekitty1296):

so it has no solution?

pooja195 (pooja195):

Wait i messed up on the first part we need to divide both sides by 4 not add once we do that we get |x+5|=4 now we can solve for x x+5=4 x+5=-4 solve both of them for x there are 2 solutions you need to figure out what they are :)

OpenStudy (zombiekitty1296):

ohhh okeii

OpenStudy (zombiekitty1296):

x=-1,x=-9??

pooja195 (pooja195):

Yep